# Machine-learned models for magnetic materials

### v1

@@ Abstract
We present a general framework for modeling materials using deep neural networks. Material represented by multidimensional characteristics (that mimic measurements) is used to train the neural autoencoder model in an unsupervised manner. The encoder is trying to predict the material parameters of a theoretical model, which is then used in a decoder part. The decoder, using the predicted parameters, reconstructs the input characteristics. The neural model is trained to capture a synthetically generated set of characteristics that can cover a broad range of material behaviors, leading to a model that can generalize on the underlying physics rather than just optimize the model parameters for a single measurement. After setting up the model we prove its usefulness in the complex problem of modeling magnetic materials in the frequency and current (out-of-linear range) domains simultaneously.
@@
materials modeling, deep neural networks, synthetic data
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Modeling different phenomena or materials is always connected with a need to tune model-specific parameters. After developing a theory that describes the underlying physics, we come to the stage where it is necessary to choose the parameters of the model, most often adjusting them to obtain a correspondence between the model characteristics and the measurements. Depending on the model complexity, the typical approach is to optimize parameters by curve fitting using least-squares or one of the vast variety of optimization algorithms, such as gradient-based methods[1], simulated annealing[2], particle swarm optimizations[3, 4], or evolutionary algorithms[5, 6]. Regardless of the algorithm chosen, optimization is always the selection of a parameter for a given set of measurement data.
With the advent of big data, deep learning, i.e. paradigm that uses deep neural networks (NNs), revolutionary breakthroughs in many conventional machine learning and pattern recognition tasks are a fact, such as computer vision or signal processing [7, 8]. The accompanied data-driven materials modeling approach [9, 10, 11, 12] also provides significant shift in materials science methodology from simple trial-and-error routine to intelligent discovery of new materials with demanding properties [13, 14, 15, 16, 17].
In this paper, we propose a similar approach but for learning any analytical model of a material in terms of intelligent fitting of its parameters. In the following, we train a machine learning model composed of a deep NN to be able to select parameters of a given theoretical model over a wide range of material characteristics at the same time, not just to single measurement data, which from machine learning perspective would be just overfilling of this single characteristic. Our goal is to build a model capable of generalizing over a variety of characteristics than a single one, which also means that the NN is able to capture the physics behind the theoretical model.
@@ II Model and data generation
To show the capability of the proposed approach, we will focus on a simple but very useful phenomenological model of magnetic materials in a wide frequency range capturing linear but also nonlinear regimes [18]. Various approaches to modeling magnetic materials in nonlinear conditions were proposed, including famous Jiles-Atherton (JA) model [19, 20], models based on Finite Element Method (FEM) that can describe local spatial dependence of magnetization [21, 22], or lumped element equivalent circuits (LEEC) suitable for direct implementation in SPICE or EMTP simulators [23, 24].
Here, we will use LEEC representing impedance characteristics of magnetic rings in both frequency and current domains, introduced in [18], and further analyzed in terms of its parameters optimization [25], also using AI methods [26]. Nanocrystalline rings are currently used in many power and electronic devices [27] to meet the requirements related to electromagnetic compatibility by suppressing electromagnetic interferences occurring in these devices and the requirements of insulation coordination by suppressing overvoltages occurring in these devices, in particular, overvoltages occurring in Gas-Insulated Switchgear. The problem of overvoltage suppression in GIS is important in the case of high-voltage GIS starting from 550 kV and ultra-high voltage GIS up to 1100 kV [28], in which fast-changing overvoltages (with frequencies of several MHz) overlap with slow-changing waveforms (with frequencies of 50/60 Hz), which are approximately constant. The significance of these conditions comes from the fact that the peak values of these overvoltages determine the design of the insulation systems of the GIS devices.
@@
To make our analysis realistic, we collect a set of measurements that will serve as a test set for the developed method. In Fig. 1(a), there is presented a family of impedances as a function of frequency $f$, taken for different bias current $I_{\mathrm{DC}}$ which controls the ring magnetization: from a linear range up to the ring saturation. The color of the measured impedance curves $Z^{\mathrm{meas}}$, that changes gradually from red, for $I_{\mathrm{DC}}=0$ A, to blue, for $I_{\mathrm{DC}}=27$ A, ($N_{\mathrm{bias}}=31$ different values in total) shows impedances for different biasings, combining this way multiple frequency characteristics controlled by the bias current $I_{\mathrm{DC}}$: $Z^{\mathrm{meas}}=Z^{\mathrm{meas}}(I_{\mathrm{DC}},f)$. Having families of characteristics, not just a single one, will be crucial when proving the model’s capability to capture broad physics rather than just fitting a single feature. Methods for measuring magnetic rings over a wide range of frequencies and saturations are not a trivial task (especially reaching the saturation limit is challenging) and are described in detail in Ref. [29].
@@
Figure 1: Impedance of a magnetic ring (a) measured (inset) as a function of frequency for various biasings that controls operating point from linear range ($I_{\mathrm{DC}}=0$ A, red curve) to the full saturation ($I_{\mathrm{DC}}=27$ A, blue curve), and (b) generated using lumped element model (inset) with continuous transition between different curves (operating regimes) and randomly selected outermost characteristics (red and blue one). The generated data will serve as a training set for the introduced neural network model, while the measurement ones will be used to test the already trained NN model.
@@
Let us now introduce the analytical model of a ring in the form of LEEC impedance synthesized out of the series of $M=7$ LR elements connected in parallel, as presented in Fig. 1(b), with the equivalent impedance in a form
@@
where $s=j2\pi{}f$, and $j$ is the imaginary unit. The goal for the NN will be to learn the model defined in Eq. 1 and predict its parameters, i.e., to find $\{L_{i},R_{i}\}$ for a given frequency characteristics over the wide range of their possible behaviors. Therefore, to train the NN model and to force it to generalize over different behaviors, we have to synthesize the training dataset using the analytical model with random parameters. Generating data and training neural models on artificially synthesized datasets with controlled properties is currently a growing trend in machine learning [30].
@@ III Neural network model
After defining the generated training and measured test datasets, let us define the NN responsible for finding the model parameters. The network will be composed of two parts: encoder and decoder – depicted in Fig. 2, and trained in an unsupervised manner. The role of the encoder is to predict the analytical model parameters $\{L_{i},R_{i}\}$ (their values define the latent space of our autoencoder), while the decoder will just calculate the frequency characteristic directly using Eq. 1 with parameters (latent space) found by the encoder.
Figure 2: Neural network basic model with an encoder-decoder structure, which is of fundamental type and used in various fields of machine learning or image processing. The encoder is made of three convolutional layers (CNN) and three fully connected (FC) layers. The CNN layers are used as feature extractors, while the FC layers are used for parameters ($\{L_{i},R_{i}\}$) predicting based on previously found representations. The decoder ($Z$) just implements an analytical formula for the impedance of the LEEC model. The goal of the network is to reconstruct the input characteristics at the output.
The goal for the NN is to reconstruct input characteristics as best as possible, with the smallest reconstruction error called loss – this means that in our training, we do not need any labels and only frequency characteristics to be reconstructed which makes our training scheme fully unsupervised.
@@
Frequency characteristics are organized on a 1D evenly spaced grid along multiple decades on the log scale ($50$ points from $135$ Hz to $10^{7}$ Hz). To build an effective neural representation, we used convolutional neural network (CNN) layers as feature extractor parts, which is a common way to process data organized on regular grids [31, 32]. Then, we applied several fully connected (FC) layers serving as parameter predictors, which are based on previously found representations (by CNN). The detailed structure of the encoder network is presented in Fig. 2: three CNN feature extractor layers are followed by three FC layers with $2M$ outputs in the last layer that predict the parameters. We select CNN as a feature extractor (backbone) part, which is common practice in signal or image processing [8, 7]; however, in principle, any other type of network could serve as backbone network at this stage, depending on applications – such as recurrent neural networks or nowadays very popular transformer networks [33] with a self-attention mechanism incorporated.
@@
$\mathrm{MSLE}\left(Z,\hat{Z}\right)=\frac{1}{N}\sum_{j=1}^{N}\log^{2}\!\left(\frac{Z_{j}}{\hat{Z}_{j}}\right).$ (3)
The autoencoder is trained over the whole training set and the sum in Eq. 2 goes over $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$ frequency characteristics $Z$ in the training set. However, in practice, during a single training step of the stochastic gradient descent (SGD) method, only a smaller random portion (mini-batch) is used, here $N_{\mathrm{batch}}=1024$. Training typically lasted $100$-$500$ epochs up to the loss saturation. To validate the NN model performance after the training, we also define the relative error of the fitting (FE):
@@ IV Basic model results
Figure 3: Results for testing the basic neural network model on the measured ring data: (a,b) examples of fitting the analytical model (solid curves), with parameters predicted by the NN model, versus measurements (dots) for the selected bias currents $I_{\mathrm{DC}}=\{0.9,6.9,12.5\leavevmode\nobreak\ \mathrm{A}\}$, (c) relative fitting error for the various bias currents $I_{\mathrm{DC}}$ plotted versus frequency, and (d) characteristic frequencies of the predicted parameters (each representing LR pair in the LEEC series) versus bias current $I_{\mathrm{DC}}$.
Let us now perform tests and apply the model to realistic, measured data as presented in Fig. 1(a). Comparison of characteristics for the analytical model (defined in Eq. 1, with parameters predicted via the NN model) and the measured one, for selected $I_{\mathrm{DC}}$ values, are presented in Fig. 3(a,b) and shows quite good fitting performance over the broad range of bias currents. It is worth recalling here that the measurement data was not a part of the training dataset. Moreover, we calculated the relative fitting error systematically, using Eq. 5 but now on the measured data, for every bias current over the frequency range as presented in Fig. 3(c). In general, we observe that the error does not exceed $30\%$, which is quite a resonable result for fitting the model over so wide ranges of frequencies, impedance amplitudes, and bias currents $I_{\mathrm{DC}}$. On average it is even smaller and equals $12.54\%$.
@@ V Siamese neural network model
The results obtained for the basic NN model are acceptable but have one substantial drawback. If we look at Fig. 3(d) which presents values of characteristic frequencies defined as $\omega_{i}=R_{i}/L_{i}$ for each of LR pairs in a series (of size $M=7$) for subsequent $I_{\mathrm{DC}}$ currents, we observe that the values are irregular and quite independent from each other which seems to be unphysical. In order to restore the more realistic behavior which is more likely to be continuous (we do not expect to have any phase transitions in the model here), the single NN model was extended to the Siamese model shown in Fig. 4.
Figure 4: Siamese neural network model composed of two copies of the basic model that share the weights and process two neighbor characteristics (that differ in bias current by $\Delta I$) at the same time enabling to train the model to predict similar (or continuously changing) parameters for similar characteristics.
The idea behind the Siamese NNs, popular in many ML areas, such as face recognition [34], is to have two (or more) copies of a network that works in parallel (but shares weights) and process different data samples at the same time. Processing multiple samples at once enables to regularize of the model behavior with respect to the similarity between data samples by adding an extra term to the loss function. In our case we can process neighbor characteristics, i.e. for $I_{\mathrm{DC}}$ and $I_{\mathrm{DC}}+\Delta I$ – see Fig. 4, and enforce parameters predicted for that data points to change continuously, by including additional loss terms in a form:
@@
in order to maintain a monotonic decrease in the parameter values with the bias current $I_{\mathrm{DC}}$, which is justified by the ring material saturation. The total monotonicity loss $\mathscr{L}_{\mathrm{mono}}(L)$ is defined in the same way as in Eq. 7. All in all, the loss function for the Siamese NN model is as follows:
@@
If we look at the results for the Siamese NN model presented in Fig. 5(a,b), where we force the model to predict similar parameters for similar characteristics, we can observe that now characteristic frequencies $\omega_{i}=L_{i}/R_{i}$ change with the bias current $I_{\mathrm{DC}}$ continuously and in some case monotonically – see Fig. 5(b). Moreover, a surprising but also very desirable property is that forcing the continuity for the change of the parameters $\{L_{i},R_{i}\}$ with $I_{\mathrm{DC}}$ at the same time has made fitting of the analytical model better than before – in Fig. 5(a) the relative fitting error typically does not exceed $15\%$ with the average value over frequencies and currents equals $7.22\%$.

### v2

# Machine-Learned Models for Power Magnetic Material Characteristics

@@ Abstract
We present a general framework for modeling power magnetic materials characteristics using deep neural networks. Magnetic materials represented by multidimensional characteristics (that mimic measurements) are used to train the neural autoencoder model in an unsupervised manner. The encoder is trying to predict the material parameters of a theoretical model, which is then used in a decoder part. The decoder, using the predicted parameters, reconstructs the input characteristics. The neural model is trained to capture a synthetically generated set of characteristics that can cover a broad range of material behaviors, leading to a model that can generalize on the underlying physics rather than just optimize the model parameters for a single measurement. After setting up the model, we prove its usefulness in the complex problem of modeling magnetic materials in the frequency and current (out-of-linear range) domains simultaneously, for which we use measured characteristics obtained for frequency up to $10$ MHz and H-field up to saturation.
@@
power magnetics, materials modeling, deep neural networks, synthetic data, magnetic ring
publicationid: pubid: $0$–$00$ 2024 IEEE
@@ I Introduction
Magnetic components, such as inductors and transformers, are vital in power electronics devices requiring dedicated modeling efforts due to specific working conditions and underlying physical principles. No fully satisfactory first-principle models have yet been proposed for power magnetic materials, and currently efforts are underway to reproduce magnetic materials characteristics using machine learning frameworks [1, 2, 3]. After developing a theory that describes the underlying physics, one comes to the stage where it is necessary to choose the parameters of the model, most often adjusting them to obtain a correspondence between the model characteristics and the measurements. Depending on the model complexity, a typical approach is to optimize parameters by curve fitting using least-squares or one of the vast variety of optimization algorithms, such as gradient-based methods[4], simulated annealing[5], particle swarm optimizations[6, 7], or evolutionary algorithms[8, 9]. Regardless of the algorithm chosen, optimization is always the selection of a parameter for a given set of measurement data.
With the advent of big data, deep learning, i.e. paradigm that uses deep neural networks (NNs), revolutionary breakthroughs in many conventional machine learning and pattern recognition tasks are a fact, such as computer vision or signal processing [10, 11]. The accompanied data-driven materials modeling approach [12, 13, 14, 15], including data-driven magnetic materials modeling [1, 2, 3], also provides significant shift in materials science methodology from simple trial-and-error routine to intelligent discovery of new materials with demanding properties [16, 17, 18, 19, 20].
In this paper, we propose a similar approach but for learning any analytical model of a material in terms of intelligent fitting of its parameters. The only requirement for the model is that it must be differentiable to allow the fitting error to be passed back (backpropagated) during the NN training. In the following, we train a machine learning model composed of a deep NN to be able to select parameters of a given theoretical model over a wide range of material characteristics at the same time, not just to single measurement data, which from machine learning perspective would be just overfilling of this single characteristic. Our aim is to construct a model capable of generalizing across various features, not just one, which also implies that the neural network can capture the physics underlying the theoretical model.
Neural networks have been applied for modeling core losses or hysteresis loops of power magnetics (e.g., nanocrystalline rings) using simple few-layer perceptrons [21, 22, 23, 24, 25, 26, 27], or recurrent NN [28]. Also, recently introduced MagNet platform [29, 30] proves that NN can be utilized to effectively learn B-H loop of magnetic material via more advanced autoencoder architectures, similar to ours.
Nanocrystalline rings are currently used in many power and electronic devices [31] to meet the requirements imposed on the devices with respect to electromagnetic compatibility by suppressing electromagnetic interferences and insulation coordination by suppressing overvoltages occurring in these devices, in particular, overvoltages occurring in Gas-Insulated Switchgear [32]).
@@ II Model and data generation
.Various approaches to modeling magnetic materials in nonlinear conditions were proposed, including famous Jiles-Atherton (JA) model [33, 34], models based on Finite Element Method (FEM) that can describe local spatial dependence of magnetization [35, 36], or lumped element equivalent circuits (LEEC) [37], suitable for direct implementation in SPICE or EMTP simulators [38, 39].
To show the capability of the proposed approach, we will focus on a phenomenological model of magnetic material in a wide frequency range capturing both linear and nonlinear H-field regimes. Here, we will use LEEC representing impedance characteristics of magnetic rings in both frequency and current domains, introduced in [40], and further analyzed in terms of its parameters optimization [41], also using AI methods [42].
@@
To make our analysis realistic, we collect a set of measurements that will serve as a test set for the developed method. In Fig. 1(a), there is presented a family of impedances as a function of frequency $f$, taken for different DC-bias current $I_{\mathrm{DC}}$ which controls the ring magnetization: from a linear range (see red curves in Fig.1a) up to the ring saturation (blue curves in Fig.1a). The color of the measured impedance curves $Z^{\mathrm{meas}}$, that changes gradually from red, for $I_{\mathrm{DC}}=0$ A, to blue, for $I_{\mathrm{DC}}=27$ A, ($N_{\mathrm{bias}}=31$ different values in total) shows impedances for different DC-bias current $I_{\mathrm{DC}}$ and frequency: $Z^{\mathrm{meas}}=Z^{\mathrm{meas}}(I_{\mathrm{DC}},f)$. Having families of characteristics, not just a single one, will be crucial when proving the model’s capability to capture broad physics rather than just fitting a single feature. Methods for measuring magnetic rings over a wide range of frequencies and DC-bias current up to saturation are not a trivial task (especially reaching the saturation limit is challenging) and are described in detail in Ref. [43].
(a) (b)
Refer to caption Refer to caption
@@
Figure 1: Impedance of a magnetic ring (a) measured as a function of frequency for various DC current biasings that controls operating point from ($I_{\mathrm{DC}}=0$ A, red curve) to the full saturation ($I_{\mathrm{DC}}=27$ A, blue curve), and (b) generated (random sample) using lumped element model with randomly selected outermost characteristics (red and blue one) and continuous transition between different curves). The generated data (b) will serve as a training set for the introduced neural network model, while the measurement ones (a) will be used to test the already trained NN model.
@@
Let us now introduce the analytical model of a ring in the form of LEEC impedance synthesized out of the series of $M=7$ L and R elements connected in parallel, as presented in Fig. 1(b), with the equivalent impedance in a form
@@
where $s=j2\pi{}f$, and $j$ is the imaginary unit. A series of parallel LR circuits was chosen as the best representation of the measurement data, whose impedance has a phase in the range of 00-$90$ degrees (indicating it is resistive-inductive). The number of $M=7$ elements was selected to ensure generality of the model and will be discussed in further sections. The goal for the NN will be to learn the model defined in Eq. 1 and predict its parameters, i.e., to find $\{L_{i},R_{i}\}$ for a given frequency characteristics over the wide range of their possible behaviors. Therefore, to train the NN model and to force it to generalize over different behaviors, we have to synthesize the training dataset using the analytical model with random parameters. Generating data and training neural models on artificially synthesized datasets with controlled properties is currently a growing trend in machine learning [44].
Figure 2: Four representative families of characteristics $Z_{j}(I_{\mathrm{DC}},f)$, where $j=1\dots N_{\mathrm{bias}}$, from the synthesized training dataset. Solid line denotes impedance amplitude and dashed line denotes impedance phase as in Fig.1b
@@ III Neural network model
Machine learning (ML) is a rapidly expanding field that has begun to revolutionize many areas, including magnetic design in power electronics [29, 30, 27, 28]. ML is a branch of AI that focuses on developing algorithms and statistical models that enable computers to perform tasks without explicit instructions. Instead, they rely on patterns learned directly from data. An autoencoder (AE) is a type of NN used to learn efficient codings of data. It is trained in an unsupervised manner to ignore irrelevant “features” and learn to represent the data in a compressed form. AEs are composed of two parts: encoder processes the input data and compresses it into a smaller, dense representation, which captures the essence of the data, while decoder, working in reverse of the encoder, takes the compressed representation and reconstructs the original input as closely as possible.
After defining the generated training and measured test datasets, let us define the NN responsible for finding the model parameters. The network will take a form of AE as illustrated in Fig. 3, and trained in an unsupervised way. The role of the encoder is to predict the analytical model parameters $\{L_{i},R_{i}\}$ (their values define the latent space of our AE), while the decoder will just calculate the frequency characteristic directly using Eq. 1 with parameters (latent space) found by the encoder.
Figure 3: Neural network basic model with an autoencoder structure, which is of fundamental type and used in various fields of machine learning or image processing. The encoder is made of three convolutional layers (CNN) and three fully connected (FC) layers. The CNN layers are used as feature extractors, while the FC layers are used for parameters ($\{L_{i},R_{i}\}$) predicting based on previously found representations. The decoder ($Z$) just implements an analytical formula for the impedance of the LEEC model. The goal of the network is to reconstruct the input characteristics at the output.
The goal for the NN is to reconstruct input characteristics as accurately as possible, minimizing the reconstruction error, defined by the loss function. This means that in our training process, we don’t require any true predictions (labels); only the frequency characteristics need to be reconstructed. This makes our training scheme fully unsupervised.
@@
Frequency characteristics are organized on a 1D evenly spaced grid along multiple decades on the log scale ($50$ points from $135$ Hz to $10^{7}$ Hz). To build an effective neural representation, we used convolutional neural network (CNN) layers as feature extractor parts, which is a common way to process data organized on regular grids [45, 46]. Then, we applied several fully connected (FC) layers serving as parameter predictors, which are based on previously found representations (by CNN). The detailed structure of the encoder network is presented in Fig. 3: three CNN feature extractor layers are followed by three FC layers with $2M$ (see. Eq.1) outputs in the last layer that predict the parameters. We select CNN as a feature extractor (backbone) part, which is common practice in signal or image processing [11, 10]; however, in principle, any other type of network could serve as backbone network at this stage, depending on applications – such as recurrent neural networks or nowadays very popular transformer networks [47] with a self-attention mechanism incorporated.
@@
$\mathrm{MSLE}\left(Z,\hat{Z}\right)=\frac{1}{N}\sum_{j=1}^{N}\left(\log(Z_{j})-\log(\hat{Z}_{j})\right)^{2}.$ (3)
The AE is trained over the whole training set and the sum in Eq. 2 goes over $N=N_{\mathrm{train}}\times N_{\mathrm{bias}}$ frequency characteristics $Z$ in the training set. However, in practice, during a single training step of the stochastic gradient descent (SGD) method, only a smaller random portion (mini-batch) is used, here $N_{\mathrm{batch}}=1024$. Training typically lasted $100$-$500$ epochs up to the loss saturation. To validate the NN model performance after the training, we also define the relative error of the fitting (FE):
@@ IV Basic model results
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 4: Results for testing the basic neural network model on the measured ring data: (a,b) examples of fitting the analytical model (solid curves), with parameters predicted by the NN model, versus measurements (dots) for the selected DC-bias currents $I_{\mathrm{DC}}=\{1.8,7.0,16.7\}\leavevmode\nobreak\ \mathrm{A}$, (c) relative fitting error for the various DC-bias currents $I_{\mathrm{DC}}$ plotted versus frequency, and (d) characteristic frequencies $f_{i}=2\pi R_{i}/L_{i}$ of the predicted NN model parameters (each line represent $f_{i}$ for one LR pair in the LEEC series) versus DC-bias current $I_{\mathrm{DC}}$.The blue color marks the frequency range where the measurements data are located. .
Let us now perform tests and apply the model to measured data as presented in Fig. 1(a). Comparison of characteristics for the analytical model (defined in Eq. 1, with parameters predicted via the NN model) and the measured one, for selected $I_{\mathrm{DC}}$ values, are presented in Fig. 4(a,b) and shows quite good fitting performance over the broad range of DC-bias currents. It is worth recalling here that the measurement data was not a part of the training dataset. Moreover, we calculated the relative fitting error systematically, using Eq. 4 but now on the measured data, for every DC-bias current over the frequency range as presented in Fig. 4(c). In general, we observe that the error does not exceed $30\%$, which is quite a resonable result for fitting the model over so wide range of frequencies, impedance amplitudes, and DC-bias currents $I_{\mathrm{DC}}$. On average it is even smaller and equals $12.54\%$.
@@ V Siamese neural network model
The results obtained for the basic NN model are acceptable but have one substantial drawback. If we look at Fig. 4(d) which presents values of characteristic frequencies defined as $\omega_{i}=R_{i}/L_{i}$ for each of LR pairs in a series (of size $M=7$) for subsequent $I_{\mathrm{DC}}$ currents, we observe that the values are irregular and quite independent from each other which seems to be unphysical. In order to restore the more realistic behavior which is more likely to be continuous (we do not expect to have any phase transitions in the model here), the standard NN model was extended to the Siamese model shown in Fig. 5.
Figure 5: Siamese neural network model composed of two copies of the basic model that share the weights and process two neighbor characteristics (that differ in DC-bias current by $\Delta I$) at the same time enabling to train the model to predict similar (or continuously changing) parameters for similar characteristics.
The idea behind the Siamese NNs, popular in many ML areas, such as face recognition [48], is to have two (or more) copies of a network that works in parallel and process different data samples at the same time. Two Siamese copies use (share) the same NN parameters (weights) while working in tandem (see Fig. 5) on two different input samples to compute two separate predictions. Throughout the training, the weights are effectively updated twice by applying distinct loss terms for each branch, along with an additional term that just compares the two predictions. Processing multiple samples at once enables to regularize of the model behavior with respect to the similarity between data samples by adding this extra term to the loss function. In our case we can process neighbor characteristics, i.e. for $I_{\mathrm{DC}}$ and $I_{\mathrm{DC}}+\Delta I$ – see Fig. 5, and enforce parameters predicted for that data points to change continuously, by including additional loss terms in a form:
@@
in order to maintain a monotonic decrease in the parameter values with the DC-bias current $I_{\mathrm{DC}}$, which is justified by the ring material saturation. The total monotonicity loss $\mathscr{L}_{\mathrm{mono}}(L)$ is defined in the same way as in Eq. 7. All in all, the loss function for the Siamese NN model is as follows:
@@
If we look at the results for the Siamese NN model presented in Fig. 6(a,b), where we force the model to predict similar parameters for similar characteristics, we can observe that now characteristic frequencies $\omega_{i}=R_{i}/L_{i}$ change with the DC-bias current $I_{\mathrm{DC}}$ continuously and in some case monotonically – see Fig. 6(b). Moreover, a surprising but also very desirable property is that forcing the continuity for the change of the parameters $\{L_{i},R_{i}\}$ with $I_{\mathrm{DC}}$ at the same time has made fitting of the analytical model better than before – in Fig. 6(a) the relative fitting error typically does not exceed $15\%$ with the average value over frequencies and currents equals $7.22\%$.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption